The mean age of 20 youth volunteers in a community project is 17.5 years with a standard deviation of 2 years. If the sample comes from an approximately normal distribution, what are the point and the interval estimates of the population mean? Use 99% confidence interval level.


1
Expert's answer
2021-05-21T15:22:02-0400

Point of estimation is the population mean age "\mu"


"\mu=\overline{x}\pm Z_{0.99}\frac{\sigma}{\sqrt{n}}=17.5\pm 2.576\cdot\frac{2}{\sqrt{20}}=17.5\pm 1.15"


"16.35\le\mu\le18.65"